OpenStudy (anonymous):

Find the distance from the point of intersection of the lines 2x+3y=10 and 3x-y=4 to the line 5x-6y=1

OpenStudy (anonymous):

First find the point set both your equations equal and solve for (x,y) (you could probably even get it by guessing) Then use the distance formula from a point to a line

OpenStudy (anonymous):

The point is 2,2 but I think I'm using the formula for the distance from a point to a line incorrectly... Could you show me how you'd do it?

OpenStudy (anonymous):

Okay, the point is (2,2) and the line is 5x-6y=1 So the formula you can use is just \[\sqrt{ax+by+c} \div \sqrt{a ^{2}}+b ^{2}\] Where the point is (x,y) and the line is ax+by+c=0

OpenStudy (anonymous):

oops.. the ax by and c terms chould all be squared under the root

OpenStudy (anonymous):

actually looking again.. they aren't squared. the first formula is right

OpenStudy (anonymous):

(without the square root)

OpenStudy (anonymous):

yeah That's the formula I have

OpenStudy (anonymous):

so the equation is actually just |ax +by+c| / sqrt(a^2 + b^2 )

OpenStudy (anonymous):

But I'm supposed to use the absolute value, right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

and that's what I think was making me do the question wrong

OpenStudy (anonymous):

so it should be 3 right?

OpenStudy (anonymous):

not -3

OpenStudy (anonymous):

Yeah, but the answer is 0.38

OpenStudy (anonymous):

So does the absolute value apply to the denominator too?

OpenStudy (anonymous):

Well both terms are squared so it won't matter for the bottom

OpenStudy (anonymous):

So... The bottom should be ... \[\sqrt{61}\] right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Hallelujah the right answer! Finally :) Thanks for your help :)

OpenStudy (anonymous):

awwyeah :)